# Alternating Perception-Reasoning for Hallucination-Resistant Video Understanding

### v1

@@ 5.4 Ablation Study
Traing Set HAVEN  [li2025vidhalluc] VideoHallucer [wang2024videohallucer]
@@ 7 Discussion
Limitations. First, the PLR paradigm relies on the precise temporal grounding capabilities of the underlying VideoLLM. While some open-source models [chen2024internvl, chen2024expanding, zhu2025internvl3, zhang2024llavanext-video] currently face challenges in timestamp generation, we anticipate that rapid advancements in foundation models (e.g., Gemini series [geminiteam2024gemini15unlockingmultimodal, comanici2025gemini]) will alleviate this bottleneck. Second, due to computational constraints, we restricted our experiments to 3B and 7B parameter scales with efficient data sizes (VideoPLR-14K and AnetHallu-117K) and controlled resolutions, rather than training larger-scale (e.g., 72B) exhaustive models. Despite these constraints, Video-PLR achieves state-of-the-art performance and high data efficiency, confirming the validity of our paradigm.

### v2

@@ 5.4 Ablation Study
Traing Set HAVEN  [14] VideoHallucer [34]
@@ 7 Discussion
Limitations. First, the PLR paradigm relies on the precise temporal grounding capabilities of the underlying VideoLLM. While some open-source models [5, 4, 49, 44] currently face challenges in timestamp generation, we anticipate that rapid advancements in foundation models (e.g., Gemini series [8, 7]) will alleviate this bottleneck. Second, due to computational constraints, we restricted our experiments to 3B and 7B parameter scales with efficient data sizes (VideoPLR-14K and AnetHallu-117K) and controlled resolutions, rather than training larger-scale (e.g., 72B) exhaustive models. Despite these constraints, Video-PLR achieves state-of-the-art performance and high data efficiency, confirming the validity of our paradigm.